# SP2T: Sparse Proxy Attention for Dual-stream Point Transformer

### v1

@@ 4.2 Ablation Study
Table 7: Ablation study about empty proxies.
@@
Table 8: Ablation study about efficiency of SP2T.
Model efficiency. Tab. 8 illustrates the model’s efficiency, evaluated in terms of accuracy and latency, for both indoor (ScanNet) and outdoor (nuScenes) datasets using a single RTX 4090. We compared our model with PTv3, PTv2, and MinkUnet. Although our model runs litter slower than PTv3, our model achieves the best balance of accuracy and latency in both indoor and outdoor datasets.
4.3 Discussion and Visualization
In this section, we discuss the limitations of the FPS sampling method and illustrate the attention map of SPA. More discussion and visualization are provided in support materials.
Over-fitting due to FPS-based Sampling. Within the ablation study, the experiments show that FPS-based sampling performs poorly compared to alternative sampling methods. To investigate this further, we visualized the point-proxy attention maps for both FPS-based and Spatial-wise sampling methods, as illustrated in Fig. 9. The visualization illustrates that the attention map resulting from FPS-based sampling exhibits static and repetitive patterns, with its attention not influenced by the proxy’s location. On the other hand, the attention map from spatial-wise sampling aligns with typical attention maps and takes the proxy location into account. Consequently, we contend that the sampling method based on FPS leads to significant overfitting of the model because of the scene leakage from FPS.
Figure 9: Visualization of the point-point attention map under FPS-based and Spatial-wise proxy sampling method. The red star represents the current point.
Figure 10: Visualization of the latency of SP2T, Sparse Proxy Attention and Global Fusion. The latency of local and global fusion include the latency of FFN.
Visualization of Latency. Fig. 10 (a) presents the latency tree map for our model. It reveals that SPA and global fusion contribute 17% to overall latency. The latency of local and global fusion include the latency of FFN. Additionally, Fig. 10 (b) & (c) display the latency tree map of each operator in SPA and global fusion. The findings indicate that SPA with global fusion significantly improves performance without imposing substantial delay.
Visualization of SPA. Fig. 11 provides a depiction of the attention map in SPA, focusing on the interaction between the point and the proxy. Specifically, Fig. 11 (a) highlights the attention map from point to proxy, and Fig. 11 (b) showcases the return path from proxy to point. Examining the visualization reveals that the SPA’s semantic mining capability is evident. In this visualization, as the chosen proxy approaches the bookshelf, both the point cloud and the proxy’s focus show a stronger affinity for the bookshelf compared to the table, showcasing the SPA’s semantic representation ability.
By processing the attention maps of point-to-proxy, proxy-to-proxy, and proxy-to-point, we can derive the corresponding point-to-point attention maps in SPA. Fig. 12 illustrates the point-to-point SPA attention map at different stages(with different downsampling rates). The visualization results indicate that SPA is capable of demonstrating semantic aggregation at different stages. This includes the focused attention of a point on one chair toward other chairs in Fig. 12 (a), the focus of a point on the wall towards the entire wall in Fig. 12 (b), and the attention of a point on the table towards other tables in Fig. 12 (c).
Figure 12: Visualization of Point to Point Attention Map of SPA of different objects and stages (with different downsampling rate). The red star represents the current point.
5 Conclusion
In this paper, we introduce SP2T, a local proxy-based dual-stream point transformer that uses sparse proxy attention to achieve global receptive field. To address the proxy sampling challenges in point clouds, we propose a spatial-wise proxy sampling method with vertex-based association, which ensures efficient and reliable proxy sampling and association calculations. Moreover, we also address the point-proxy cross-attention problem by incorporating sparse proxy attention with table-based relative bias, which enhances precise interactions between proxy and point features. Comprehensive experiments across multiple datasets show that our model achieves state-of-the-art performance in various downstream tasks. For limitations, we claim that the implementation of SPA and network parameters can be further refined. Compared to PTV3, our method uses more time and memory, and we will continue this research to address these challenges.
Appendix A Experiment and Discussion
Indoor Sem. Seg. ScanNet [10] ScanNet200 [32]
Methods with same TTA Val Test Val Test
MinkUnet [7] 72.2 73.4 25.0 25.3
OctFormer [40] 74.6 70.7 31.9 31.0
Swin3D [45] 76.6 71.4 - -
PTv3 [44] 77.5 73.6 35.2 34.0
SP2T 78.7 74.9 37.0 35.2
Table 9: Indoor instance segmentation with same TTA between Val and Test set. Reproduction using code of Pointcept. Official weight and code.
Indoor Sem. Seg. ScanNet [10] ScanNet200 [32]
Methods Val Test Val Test
MinkUnet [7] 72.2 73.4 25.0 25.3
SP2T w/ MinkUnet 74.7($\uparrow$2.5) - 28.3($\uparrow$3.3) -
PTv3 [44] 77.5 73.6 35.2 34.0
Table 10: Transfer study of local fusion.
A.1 Result Gap in Test Set
According to [33, 45, 44], there is a large test time augmentation (TTA) difference between the val and test sets of Scannet [10] and Scannet200 [32]. The additional TTA includes incorporating data from the validation set for training, combining results from multiple training models [20], and applying over-segmentation [13].
For fair comparison, we evaluated some SOTA models [7, 40, 45, 44] with an official or reproduction weights file using Val’s TTA, as shown in Fig. 9. The experimental findings indicate that, without employing additional TTA, the result in the test set for the majority models [40, 45, 44] tends to be less than those on the validation set, rather than exceeding them. And it can be found that our model archive the best result both in the val and test set of Scannet and Scannet200.
We are still working on over-segmentation and may update our model’s test results employing over-segmentation in the final version of the paper. Furthermore, we recommend that future research ensure consistency in the TTA between the validation set and the test set, or at least make the TTA on the test set openly available. In the field of 3D understanding, the focus should be on improving network design and training methodologies rather than on using more testing tricks.
A.2 Transfer Study of Local Fusion
We also use MinkUnet [7] as the local fusion module to test the portability of our design. Tab. 10 shows the performance of MinkUnet [7] and PTv3 [44] and its variants with our proposed sparse proxy attention mechanism on the Scannet [10] and Scannet200 [32] datasets. The experimental results demonstrate that our network structure is plug-and-play and can integrate seamlessly with various local fusion networks, thereby enhancing the receptive field and boosting the network’s performance.
A.3 Visualization
Figure 13: Result Comparison in ScanNet.
Figure 14: Failed cases in ScanNet.
Failed Cases. Fig. 14 illustrates the failed cases of our method on the Scannet dataset [10]. In Fig. 14, it is apparent that the proxy is failed to address the classification error attributed to location fusion. Consequently, achieving a proper balance between local and global information still requires further investigation.
A.4 Open Source.
The supplementary material includes the training and testing code along with the configuration files. In addition, we have shared a link to an anonymous GitHub repository which contains the training logs and the val / test result for our model. Furthermore, we provide test results for other models including Swin3d [45], OctFormer [40], and PTv3 [44] in the repository for both Scannet and Scannet200 for peer review. We will release the weight of our model if the paper is accepted.
@@ A.3 Visualization
optimizer AdamW datasets NuScenes
@@
criteria CrossEntropy (1) batch size 12
Lovasz [2] (1) datasets NuScenes
@@
Appendix B Implementation Details
@@
B.1 Training Settings
Datasets and metrics. The ScanNet dataset [10, 32], frequently utilized in indoor real-world down-stream tasks, contains 1,513 room scans derived from RGB-D frames, with 1,201 scenes designated for training and 312 reserved for validation. Each point was categorized into one of the 20 semantic labels in ScanNet [10] and 200 semantic labels in ScanNet200 [32]. In contrast, the S3DIS dataset [1] covers 271 rooms in six areas within three buildings with 13 categories. nuScenes [3] consists of 40,157 annotated samples, each containing six monocular camera images that cover a 360-degree field of view and a 32-beam LiDAR. According to the specifications of nuScenes, the dataset is composed of 1000 scenarios, 1.4M images, and 400K point clouds. The training set covers 700 scenarios, and both the validation and test sets contain 150 scenarios each. For metrics, we utilize the mean class-wise intersection over union (mIoU) as the principal metric in ScanNet, ScanNet200, and S3DIS. Furthermore, following previous work, area 5 in S3DIS is designated for testing with a 6-fold cross-validation.
Outdoor semantic segmentation. Similarly to indoor segmentation, Tab. 12 outlines the training parameters for SP2T when applied to outdoor segmentation. Similarly to our approach for indoor segmentation, the model undergoes a two-stage training process. Initially, in the first stage, a distinct local fusion network is trained specifically for nuScenes. Then, for the second stage, the model is initialized with the parameters of this local fusion network and continues to train.
B.2 Model Settings

### v2

@@ 4.2 Ablation Study
[…]
Table 6: Ablation study about hyper-parameter of SP2T in ScanNet. The baseline structure of these models is Tab. 5 #10. Asso. means the number of associations. Wise. means Spatial-wise sampling. E.P. means Empty Proxies. OOM means out of memory.
Indoor Sem. Seg. ScanNet [11] ScanNet200 [43]
Methods Val Test Val Test
MinkUnet [9] 72.2 73.4 25.0 25.3
SP2T w/ MinkUnet 75.7($\uparrow$3.5) - 30.3($\uparrow$5.3) -
PTv3 [56] 77.5 73.6 35.2 34.0
Table 7: Transfer study of local fusion.
Proxy Sampling Method. Tab. 6 rows 11 to 14 provides the results of the ablation study on proxy sampling methods. The experiment indicates that the spatial-wise sampling method yields the best model accuracy among various sampling methods. The Fix Number and Fix Size-based methods struggle with geometrically diverse point regarding the FPS-based methods, reducing accuracy and failure training. In contrast, the spatial-wise sampling approach can flexibly determine the proxy spacing and improve performance. The more discussion and visualization is provided in supplementary materials regarding the FPS-based method.
Empty Proxies. Unlike image, grid sampling in a point cloud unavoidably results in empty proxies, which lack any point cloud association. Tab. 6 rows 15 to 16 compares SP2T accuracy with and without considering empty proxies during global fusion. In the experiment, empty proxies are initialized with zero features and participate in global fusion to encode spatial voids. The experiment indicates that empty proxies enhance the accuracy of the model. The experiment proves that ‘nothing’ is also critical information in the point cloud. Although empty proxies lack points, the absence itself conveys spatial structure information, thereby contributing to an understanding of overall spatial features for points.
Transfer Study of Local Fusion. We also use MinkUnet [9] as the local fusion module to test the portability of our design. Tab. 7 shows the performance of MinkUnet [9] and PTv3 [56] and its variants with our proposed sparse proxy attention mechanism on the Scannet [11] and Scannet200 [43] datasets. The experimental results demonstrate that our network structure is plug-and-play and can integrate seamlessly with various local fusion networks, enhancing the receptive field and boosting the network’s performance.
4.3 Visualization
In this section, we illustrate the attention map and latency of SPA. More visualization and failed cases are provided in supplementary materials.
Figure 7: Visualization of the latency of SP2T, Sparse Proxy Attention, and Global Fusion. The latency of local and global fusion includes the latency of FFN.
Visualization of Latency. Fig. 7 (a) presents the latency tree map for our model. It reveals that SPA and global fusion contribute 17% to overall latency. The latency of local and global fusion includes the latency of FFN. Additionally, Fig. 7 (b) & (c) display the latency tree map of each operator in SPA and global fusion. The findings indicate that SPA with global fusion significantly improves performance without imposing substantial delay.
Visualization of SPA. Fig. 8 depicts the attention map in SPA, focusing on the interaction between the point and the proxy. Specifically, Fig. 8 (a) highlights the attention map from point to proxy, and Fig. 8 (b) showcases the return path from proxy to point. Examining the visualization reveals that the SPA’s semantic mining capability is evident. In this visualization, as the chosen proxy approaches the bookshelf, the point cloud and the proxy’s focus show a stronger affinity for the bookshelf than the table, showcasing the SPA’s semantic representation ability.
By processing the attention maps from point-to-proxy, proxy-to-proxy, and proxy-to-point, Fig. 9 illustrates the point-to-point SPA attention map at different stages(with different downsampling rates). Specifically, Fig. 9 (a), (b), and (c) illustrate the visualization results corresponding to 1-st, 2-rd, and 3-th SP2TLayer, respectively. It is noted that both object semantics and TRB together influence the model’s attention. Initially, in the early stage of the model, TRB primarily directs the attention map by emphasizing the spatial pattern. Object semantics increasingly dominate the model’s attention as the model progresses to a deeper stage.
Figure 9: Visualization of point-to-point attention map of SPA of different objects and stages. The red star represents the current point. Specifically, (a), (b), and (c) illustrate the visualization results corresponding to 1-st, 2-rd, and 3-th SP2TLayer.
5 Conclusion
We present a novel method called Sparse Proxy Point Transformer (SP2T) designed for 3D understanding. It incorporates a local proxy-based dual-stream transformer and uses sparse proxy attention to enable a global receptive field. Key innovations include a spatial-wise sample method for robust sampling, sparse proxy attention for efficient proxy-point feature interaction, and a dual-stream structure to balance local and global information processing. SP2T demonstrates state-of-the-art performance across various tasks despite being improvable in terms of speed. Future improvements focus on parameter tuning and optimizing dual-stream dimensional balance.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62433003) and the National Natural Science Foundation of China (62476017).
Appendix A Experiment and Discussion
Indoor Sem. Seg. ScanNet [11] ScanNet200 [43]
Methods with same TTA Val Test Val Test
MinkUnet [9] 72.2 73.4 25.0 25.3
OctFormer [52] (Rep.) 74.6 70.7 31.9 31.0
OctFormer [52] (Off.) 75.7 - 32.6 -
Swin3D [59] (Off.) 76.4 - - -
Swin3D [59] (Rep.) 76.6 71.4 - -
PTv3 [56] (Off.) 77.5 73.6 35.2 34.0
SP2T 78.7 74.9 37.0 35.2
Table 8: Indoor instance segmentation with same TTA between Val and Test set. Rep. means the model uses the code of Pointcept and reproduces it by ours. Off. means the model uses official weight and code.
Algorithm 1 Pytorch-Style Pesade-code of Spatial-wise Sampling
def Spatial_Wise_Sampling(
s_min: float, s_max: float, # min/max cell size
cnt_low: int, cnt_high: int, # target count range
max_iter: int, # max iterations
grid_range: float # AABB size
) -> float: # optimal grid size
l, r = s_min, s_max
for _ in range(max_iter):
grid_size = (l + r) / 2
grid_shape = ceil(grid_range / grid_size)
cell_count = prod(grid_shape)
if cnt_low <= cell_count <= cnt_high:
return grid_size
elif cell_count < cnt_low:
r = grid_size # too sparse
else:
l = grid_size # too dense
return (l + r) / 2 # fallback
A.1 ScanNet Test Set
According to [44, 59, 56], there is a significant test time augmentation (TTA) difference between the val and test sets of Scannet [11] and Scannet200 [43]. The additional TTA includes incorporating data from the validation set for training, combining results from multiple training models [24], and applying over-segmentation [14].
For a fair comparison, we evaluated some SOTA models [9, 52, 59, 56] with an official or reproduction weights file using Val’s TTA, as shown in Fig. 8. The experimental findings indicate that, without employing additional TTA, the result in the test set for the majority models [52, 59, 56] tends to be less than those on the validation set, rather than exceeding them. And it can be found that our model archive the best result both in the val and test set of Scannet and Scannet200.
We are still working on over-segmentation and may update our model’s test results employing over-segmentation in the final version of the paper. Furthermore, we recommend that future research ensure consistency in the TTA between the validation and test sets or at least make the TTA on the test set openly available. In 3D understanding, the focus should be on improving network design and training methodologies rather than using more testing tricks.
A.2 Pesade-code of Spatial-wise Sampling
The pesade-code of spatial-wise sampling is shown in Alg. 1. Spatial-wise sampling efficiently discerns the ideal proxy spacing by considering the AABB sizes of various points. The method is designed to maintain the proxies count within the bounds of $N_{max}$ and $N_{min}$, using a bisection approach to determine the optimal proxy spacing $L_{p}$. If the number of proxies exceeds $N_{max}$, the proxy spacing $L_{p}$ is reduced, and the total number of proxies is recalculated.
A.3 Visualization
Figure 10: Result Comparison in ScanNet.
Figure 11: Failed cases in ScanNet.
Failed Cases. Fig. 11 illustrates the failed cases of our method on the Scannet dataset [11]. In Fig. 11, it is apparent that the proxy has failed to address the classification error attributed to location fusion. Consequently, achieving a proper balance between local and global information still requires further investigation.
Figure 12: Visualization of the point-point attention map under FPS-based sampling. The red star represents the current point.
Over-fitting due to FPS-based Sampling. Within the ablation study, the experiments show that FPS-based sampling performs poorly compared to alternative sampling methods. We visualized the point-proxy attention maps for FPS-based sampling methods to investigate this further, as illustrated in Fig. 12. The visualization demonstrates that the attention map resulting from FPS-based sampling exhibits static and repetitive patterns, with its attention not influenced by the proxy’s location. Consequently, we contend that the sampling method based on FPS leads to significant overfitting of the model because of the scene leakage from FPS.
TRB Share TRB mIoU mAcc allAcc Time
[…]
Table 9: Ablation study about sharing of TRB.
Figure 13: (a) MSE of TRB in different proxy-point interaction layers. (b) MSE of TRB in different proxy-proxy interaction layers.
A.4 Shared Table Relative Bias
During the implementation phase, the point and proxy positions remain unchanged within the same layer. This means that for each instance of sparse attention in this layer, the relative position input provided to the relative position encoding module remains constant. Consequently, it is possible to compute the relative bias for a layer with a single invocation. Building on this optimization, we share all relative bias values across each layer, thus reducing model complexity and computational demand. Additionally, since the proxy positions are constant throughout the network, we test sharing the relative bias amongst proxies across the entire network and different layers.
Tab. 9 compares the accuracy of the model w/ and w/o the shared TRB. The experiment shows that the shared TRB improves the model’s accuracy and reduces the inference time. Furthermore, Fig. 13 shows the MSE distance for TRB during point-proxy and proxy-proxy interaction in different layers. TRB demonstrates stage-level similarity in the proxy-proxy interaction, while all TRB is similar in the point-proxy interaction. Consequently, the sharing of TRB improves model accuracy and reduces inference time.
@@ A.4 Shared Table Relative Bias
Table 10: Ablation study about efficiency of SP2T. PTv3 refers to PTv3 with an increased number of channels to maintain the same latency as SP2T.
A.5 Model efficiency
Tab. 10 presents the model’s performance based on accuracy and latency for both the indoor (ScanNet) and outdoor (nuScenes) datasets, tested on a single RTX 4090. It is important to note that scaling the PTv3 model to match the size of SP2T does not substantially improve the metrics for either dataset. Our model achieves an ideal compromise between accuracy and speed while maintaining consistent accuracy.
Appendix B Implementation Details
B.1 Training Settings
@@
Datasets and metrics. The ScanNet dataset [11, 43], frequently utilized in indoor real-world down-stream tasks, contains 1,513 room scans derived from RGB-D frames, with 1,201 scenes designated for training and 312 reserved for validation. Each point was categorized into one of the 20 semantic labels in ScanNet [11] and 200 semantic labels in ScanNet200 [43]. In contrast, the S3DIS dataset [1] covers 271 rooms in six areas within three buildings with 13 categories.
nuScenes [4] consists of 40,157 annotated samples, each containing six monocular camera images that cover a 360-degree field of view and a 32-beam LiDAR. According to the specifications of nuScenes, the dataset comprises 1000 scenarios, 1.4M images, and 400K point clouds. The training set covers 700 scenarios, and the validation and test sets contain 150 scenarios each. SemanticKITTI [2] originates from the KITTI Vision Benchmark Suite and is comprised of 22 sequences, with 19 designated for training and the other 3 reserved for testing. Waymo [48] is a frequently utilized benchmark for outdoor 3D perception, comprising a total of 1,150 point cloud sequences (exceeding 200K frames). Each frame encompasses an extensive perception range of 150m × 150m.
For segmentation metrics, we utilize the mean class-wise intersection over union (mIoU) as the principal metric in ScanNet, ScanNet200, and S3DIS. Furthermore, following previous work, area 5 in S3DIS is designated for testing with a 6-fold cross-validation. For dectection metircs, all results are assessed by the conventional protocol employing 3D mean Average Precision (mAP) and its weighted version based on heading accuracy (mAPH).
@@
optimizer AdamW batch size 12
@@
criteria CrossEntropy (1) datasets NuScenes
Lovasz [3] (1) Sem.KITTI
Waymo
@@
Outdoor Detection
Config Value Config Value
optimizer AdamW datasets Waymo
scheduler Cosine weight decay 1e-2
framework CenterPoint batch size 12
First Stage:
learning rate 3e-3 epochs 24
block lr scaler 1e-1 warmup epochs 0
Second Stage:
learning rate 3e-4 epochs 12
block lr scaler 1.0 warmup epochs 0
Table 13: Outdoor object detection settings.
Outdoor semantic segmentation. Similarly to indoor segmentation, Tab. 12 outlines the training parameters for SP2T when applied to outdoor segmentation. Similarly to our approach for indoor segmentation, the model undergoes a two-stage training process. In the first stage, a distinct local fusion network is explicitly trained for dataset [4, 2]. Then, for the second stage, the model is initialized with the parameters of this local fusion network and continues to train.
Outdoor object detection. Tab. 13 outlines the training parameters for SP2T when applied to outdoor object detection. The model is also trained through a two-stage approach. Initially, a specific local fusion network is exclusively trained for Waymo [48]. Subsequently, in the second stage, the model starts with the parameters from this local fusion network, and training is resumed.
B.2 Model Settings